### v1

@@ Abstract
Cellular traffic forecasting is a critical task that enables network operators to efficiently allocate resources and address anomalies in rapidly evolving environments. The exponential growth of data collected from base stations poses significant challenges to processing and analysis. While machine learning (ML) algorithms have emerged as powerful tools for handling these large datasets and providing accurate predictions, their environmental impact, particularly in terms of energy consumption, is often overlooked in favor of their predictive capabilities.
This study investigates the potential of two bio-inspired models: Spiking Neural Networks (SNNs) and Reservoir Computing through Echo State Networks (ESNs) for cellular traffic forecasting. The evaluation focuses on both their predictive performance and energy efficiency. These models are implemented in both centralized and federated settings to analyze their effectiveness and energy consumption in decentralized systems. Additionally, we compare bio-inspired models with traditional architectures, such as Convolutional Neural Networks (CNNs) and Multi-Layer Perceptrons (MLPs), to provide a comprehensive evaluation. Using data collected from three diverse locations in Barcelona, Spain, we examine the trade-offs between predictive accuracy and energy demands across these approaches.
The results indicate that bio-inspired models, such as SNNs and ESNs, can achieve significant energy savings while maintaining predictive accuracy comparable to traditional architectures. Furthermore, federated implementations were tested to evaluate their energy efficiency in decentralized settings compared to centralized systems, particularly in combination with bio-inspired models. These findings offer valuable insights into the potential of bio-inspired models for sustainable and privacy-preserving cellular traffic forecasting.
@@ 1 Introduction
This study is motivated by the rapid growth of data generated by telecommunication systems, which is driven by the increasing use of connected devices. The large amounts of data involved require sophisticated predictive models that need significant computational resources to process. As a result, the training of these models often consumes substantial amounts of energy, time, and storage, creating challenges for both efficiency and sustainability.
One major concern is the environmental impact of training such resource-intensive models. As demand for data-driven solutions grows, the energy consumption and carbon emissions from these processes increase. This highlights the need for approaches that can balance strong predictive performance with lower energy consumption.
Figure 1: A architecture for real-time cellular traffic forecasting  [3].
A source of inspiration for this work is the efficiency of the human brain, which processes information using far less energy than modern machine learning systems. For example, training a model like GPT-3 is estimated to consume around 190,000 kWh of energy, whereas the human brain operates on just 12–20 W [4]. By mimicking the brain’s processes, bioinspired models like Spiking Neural Networks (SNNs) offer a promising way to make machine learning more energy-efficient.
In this study, we explore the use of bioinspired SNNs and Reservoir Computing (RC) with Echo State Networks (ESNs) for cellular traffic forecasting. We examine several types of Leaky Integrate-and-Fire (LIF) neurons in SNNs, including Lapicque, Leaky, RLeaky, Synaptic, and Alpha neurons [4], as well as ESNs [5]. These models are compared with traditional machine learning methods, such as Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs). Both centralized and federated learning settings are used in the evaluation to account for modern distributed computation approaches. Federated learning, in particular, reduces the need for centralized data storage, offering a more energy-efficient way to train models.
This study evaluates the trade-offs between predictive accuracy and energy efficiency. While energy-efficient models often show reduced prediction performance, this work aims to explore ways to strike a balance between sustainability and accuracy. To achieve this, we use a Sustainability Index [6] that combines key metrics like power consumption, prediction error, and dataset size to provide a comprehensive evaluation of model efficiency.
@@
The results aim to guide researchers and practitioners toward developing machine learning models that balance high performance with lower energy use. The rest of the paper is structured as follows: Section 2 re/views related work on Spiking Neural Networks (SNNs), Reservoir Computing (RC), and Federated Learning. Section 3 describes the dataset and research problem, along with the models and evaluation methods. Section 4 explains the experimental setup and methodology. Section 5 presents the results and analysis. Section 6 discusses the findings, challenges, and future directions. Finally, Section 7 concludes by summarizing the key outcomes and contributions of the study.
@@ 2.1 Traffic Prediction
Lee et al. (2024)  [12] explored federated learning applications beyond traffic forecasting, focusing on comprehensive mobile network management. Their study emphasized the integration of federated models into resource allocation and small-cell base station management, showcasing the broader impact of this technology on network efficiency.
@@
2.2 SNN in Traffic Prediction
Recent advancements in spiking neural networks (SNNs) demonstrate their potential for efficient and accurate network traffic prediction, particularly in energy-constrained environments. In our previous work  [3] we explored SNNs in the context of cellular traffic forecasting, emphasizing their ability to reduce power consumption compared to conventional machine learning models. By leveraging the bio-inspired architecture of SNNs, the study addressed the growing demand for environmentally sustainable machine learning solutions in telecommunications. We highlighted how SNNs effectively capture temporal dynamics in network traffic patterns, enabling precise predictions with significantly reduced energy footprints.
Similarly, Kang et al. (2024)  [16] investigated the use of SNNs in neuromorphic systems for traffic hotspot prediction in network-on-chip (NoC) architectures. Their study focused on utilizing liquid state machines (LSMs)—a specific form of reservoir computing based on spiking dynamics—to predict and mitigate traffic congestion hotspots in NoC environments. The findings showcased how SNN-based approaches not only improve computational efficiency but also enhance routing strategies, contributing to a more energy-efficient processing pipeline.
@@ 2.3 RC in Traffic Prediction
Further advancing the field, Zhang and Vargas (2023)  [18] applied reservoir computing in the context of 5G and beyond networks. Their study emphasized how reservoir computing architectures enhance forecasting accuracy while addressing sustainability challenges in mobile networks. By leveraging innovative learning techniques, their work provides actionable insights for integrating reservoir computing frameworks into modern telecommunications infrastructures, ensuring both predictive performance and energy efficiency.
Collectively, these studies underline reservoir computing’s transformative potential for cellular traffic forecasting, particularly in energy-constrained and dynamic environments like 5G networks. Its ability to balance computational efficiency with prediction accuracy makes it a critical tool for future network management and optimization.
@@ 2.4 Federated Bio Inspired in Traffic Prediction
aOur study distinguishes itself by combining SNNs and RC models within federated learning environments to assess their predictive capacity and energy efficiency. Unlike prior studies that prioritize either centralized configurations or accuracy alone, we employ a Sustainability Index (See Section 3) to evaluate the trade-offs across multiple dimensions, including power consumption, prediction error, and data size. This comprehensive approach, coupled with comparative evaluations against traditional architectures such as MLPs and CNNs, provides deeper insights into how bioinspired models can support sustainable and scalable solutions for cellular traffic forecasting. By leveraging real-world datasets and consistent experimental setups, our work offers a holistic framework for advancing energy-efficient and privacy-preserving network management practices.
@@ 3.1 Dataset
As showed in [8], the dataset presents non Independent and Identically Distributed (non-IID) distibution, with the presence of three different skews, i.e. quantity, quality and temporality. This differentiation among respective BS is depicted in Fig. 3.1, indicating a challenging ML task.
@@ 3.3 ML Models
  • Lapicque Neuron (Original LIF Neuron): The Lapicque neuron implements the fundamental circuit approach, originating from biology. Ion channels of a physical neuron’s membrane are modeled with a resistor and the membrane itself is modeled with a capacitor. We check the potential ($U$) in the capacitor, and if this potential reaches a predefined threshold, the neuron fires an output spike. We also choose a mechanism that is activated when a spike is fired to reset our neuron. Thus, the tunable parameters in this approach are the resistance ($R$), capacitance ($C$), threshold, the number of spikes for each input ($I_{in}$) and the reset mechanism. Formally, Lapicque neuron [4, 21] is modeled through the following function:
@@
    When the resulting value of $U$ is higher than the specified threshold, the neuron will fire a spike.
  • Leaky Neuron: The initial description of the LIF model, as described in [4] and [21], is distinguished by a broad set of hyperparameters. As the model is scaled to support a universal SNN architecture, managing these parameters becomes challenging. For this type of neuron, we use the following equation:
@@
  • Recurrent Leaky (RLeaky) Neuron: This implementation, as discussed in [4, 22], employs a logic similar to that of Leaky neuron, yet it uniquely scales the output spikes of a neuron by a factor of $V$ and reincorporates them into the input. Specifically, spikes produced by a layer of spiking neurons are forwarded to the following fully connected (or convolutional) layer, where they are combined with the incoming input before being processed by the spiking neuron model.
  • Synaptic Based Neuron: In the neuron models discussed previously, the premise has been that a voltage spike upon input instantaneously elevates the synaptic current, subsequently influencing the membrane potential. However, in practical scenarios, a spike initiates a progressive discharge of neurotransmitters from the pre-synaptic to the post-synaptic neuron. The synaptic conductance-based [23, 21] LIF model accommodates the slow temporal progression of the input current. Thus, the equation is transformed to:
@@
  • Reservoir Computing Echo State: Reservoir computing leverages a fixed, randomly initialized dynamic reservoir to project input data into a high-dimensional feature space. The Echo State Network (ESN) is a prominent implementation of this paradigm, particularly suited for sequential data processing. The reservoir consists of recurrently connected nodes, which transform the input time-series into a rich set of temporal features. Unlike traditional neural networks, the reservoir weights remain fixed, and only the output layer is trained, significantly reducing the computational complexity. In our implementation, the input data $(S,d)$ is mapped to the reservoir with a dimensionality of $N_{r}$ nodes. The output is then computed as a linear combination of the reservoir states and passed to a fully connected layer for final predictions. This architecture facilitates efficient processing of multivariate time-series with minimal parameter tuning and low energy consumption.
@@ 3.4 Evaluation Metrics
The primary metric used to quantify prediction error for the time-series forecasting task is the normalized root mean squared error ($\mathrm{NRMSE}$), defined as:
$\mathrm{NRMSE^{(k)}}=\frac{1}{\overline{x}}\sqrt{\frac{\sum_{t=1}^{T}\left(\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}\right)}{Td^{\prime}}},$ (5)
where $\overline{x}$ represents the mean of the target observations, computed as:
$\overline{x}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\left(\sum_{i=1}^{d^{\prime}}x_{t,i}\right),$ (6)
and $\hat{x}_{t,i}$ and $x_{t,i}$ denote the predicted and actual values at time $t$ for feature $i$, $T$ is the total number of time steps, and $d^{\prime}$ is the number of features.
In addition to $\mathrm{NRMSE}$, we also employ other common error metrics to evaluate model performance:
  • Mean Absolute Error (MAE):
    $\mathrm{MAE^{(k)}}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left|\hat{x}_{t,i}-x_{t,i}\right|,$ (7)
  • Root Mean Square Error (RMSE):
    $\mathrm{RMSE^{(k)}}=\sqrt{\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}},$ (8)
  • Mean Square Error (MSE):
    $\mathrm{MSE^{(k)}}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}.$ (9)
Beside the accuracy of the implemented algorithms, we also utilize the Sustainability Indicator $\mathcal{S}$ introduced in [6] to assess the resulting trade-off between the prediction error and the energy consumption. $\mathcal{S}$ enables a fair comparison among different model architectures and can be calculates using the following formula:
@@
where $E_{\text{Val}}$ is the validation error (in this work, we consider NMRSE), $C_{\text{Tr}}$ represents the total energy consumed for model training in Watt Hours (Wh) and $D$ is the data size to be transmitted in kB.
@@ 4 Experimental Setup
To ensure the integrity and validity of our comparative analysis, we maintained consistency across all experimental settings. The architecture of the SNN models was standardized to include a single hidden layer with 128 neurons and five output neurons corresponding to the target variables.
@@
In the output layer, however, neurons lack a reset mechanism—a deliberate design choice to better address the requirements of the time-series forecasting task. This configuration ensures that the predicted output corresponds directly to the membrane potential, avoiding distortions caused by resets.
Additionally, input spikes are encoded directly from the input data at specific timesteps, enabling the network to effectively convert non-spiking data into a format compatible with spiking mechanisms. This approach enhances the model’s ability to process time-series data within the spiking neural network framework.
The size of the dataset used for training the model is also a crucial factor. Training on a smaller dataset inevitably results in lower power consumption compared to a larger one. This relationship is captured by the Sustainability Index (S), which includes dataset size as a parameter. To ensure consistency, we used the same dataset and size across all experiments.
For implementation, we utilized the PyTorch library [25] for traditional models such as MLPs and CNNs. The SNN models were implemented using the snnTorch library [4], which is tailored for gradient-based learning and provides pre-designed spiking neuron models that seamlessly integrate as recurrent activation units. Energy consumption during the training process was measured using the CarbonTracker Python module [26]. CatbonTracker is a tool designed to estimate and monitor the carbon emissions associated with computational processes, helping organizations and researchers assess and reduce their environmental impact.
Centralized training was conducted over 150 epochs with a batch size of 128. In the federated learning experiments, we used 10 federated steps with 3 local epochs per client with a central server, simulating the setup to maintain consistency.
The experiments were conducted on a Windows 11 workstation equipped with an AMD Ryzen 7 6800HS CPU and 16 GB of memory.

### v2

@@ Abstract
Cellular traffic forecasting is a critical task that enables network operators to efficiently allocate resources and address anomalies in rapidly evolving environments. The exponential growth of data collected from base stations poses significant challenges to processing and analysis. While machine learning (ML) algorithms have emerged as powerful tools for handling these large datasets and providing accurate predictions, their environmental impact, particularly in terms of energy consumption, is often overlooked in favor of their predictive capabilities. This study investigates the potential of two bio-inspired models: Spiking Neural Networks (SNNs) and Reservoir Computing through Echo State Networks (ESNs) for cellular traffic forecasting. The evaluation focuses on both their predictive performance and energy efficiency. These models are implemented in both centralized and federated settings to analyze their effectiveness and energy consumption in decentralized systems. Additionally, we compare bio-inspired models with traditional architectures, such as Convolutional Neural Networks (CNNs) and Multi-Layer Perceptrons (MLPs), to provide a comprehensive evaluation. Using data collected from three diverse locations in Barcelona, Spain, we examine the trade-offs between predictive accuracy and energy demands across these approaches. The results indicate that bio-inspired models, such as SNNs and ESNs, can achieve significant energy savings while maintaining predictive accuracy comparable to traditional architectures. Furthermore, federated implementations were tested to evaluate their energy efficiency in decentralized settings compared to centralized systems, particularly in combination with bio-inspired models. These findings offer valuable insights into the potential of bio-inspired models for sustainable and privacy-preserving cellular traffic forecasting.
@@ 1 Introduction
This study is motivated by the rapid growth of data generated by telecommunication systems, which is driven by the increasing use of connected devices. The large amounts of data involved require sophisticated predictive models that need significant computational resources to process. As a result, the training of these models often consumes substantial amounts of energy, time, and storage, creating challenges for both efficiency and sustainability. This underscores the need for approaches that balance strong predictive performance with reduced energy usage. This consideration is particularly important in the context of federated learning, where models are trained on edge devices with limited computational and energy resources. In such scenarios, it is critical to identify models that can be effectively trained under constrained conditions.
Furthermore, the availability of models with low power and computational requirements is essential for ensuring service continuity during emergencies—such as power outages or infrastructure disruptions—where maintaining operation at reduced capacity is vital.
In this work, we focus on identifying and analyzing the trade-offs between predictive performance and power consumption across different model architectures and deployment settings.
In this study, we explore the use of bioinspired SNNs and Reservoir Computing (RC) with Echo State Networks (ESNs) for cellular traffic forecasting. We examine several types of Leaky Integrate-and-Fire (LIF) neurons in SNNs, including Lapicque, Leaky, RLeaky, Synaptic, and Alpha neurons [3], as well as ESNs [4]. These models are compared with traditional machine learning methods, such as Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs). Both centralized and federated learning settings are used in the evaluation to account for modern distributed computation approaches. Federated learning, in particular, reduces the need for centralized data storage, offering a more energy-efficient way to train models and we aim to evaluate the trade-offs between predictive accuracy and energy efficiency. To achieve this, we use a Sustainability Index that combines key metrics like power consumption, prediction error, and a size parameter to provide a comprehensive evaluation of model efficiency.
@@
The results aim to guide researchers and practitioners toward developing machine learning models that balance high performance with lower energy use. The rest of the paper is structured as follows: Section 2 reviews related work on Spiking Neural Networks (SNNs), Reservoir Computing (RC), and Federated Learning. Section 3 describes the dataset and research problem, along with the models and evaluation methods. Section 4 explains the experimental setup and methodology. Section 5 presents the results and analysis. Section 6 discusses the findings, challenges, and future directions. Finally, Section 7 concludes by summarizing the key outcomes and contributions of the study.
@@ 2.1 Traffic Prediction
2.2 SNN in Prediction Tasks
Recent advancements in spiking neural networks (SNNs) demonstrate their potential for efficient and accurate network traffic prediction, particularly in energy-constrained environments. In our previous work  [13] we explored SNNs in the context of cellular traffic forecasting, emphasizing their ability to reduce power consumption compared to conventional machine learning models. By leveraging the bio-inspired architecture of SNNs, the study addressed the growing demand for environmentally sustainable machine learning solutions in telecommunications.
Similarly, Kang et al. (2024)  [14] investigated the use of SNNs in neuromorphic systems for traffic hotspot prediction in network-on-chip (NoC) architectures. The findings showcased how SNN-based approaches not only improve computational efficiency but also enhance routing strategies, contributing to a more energy-efficient processing pipeline.
Early work by Ghosh-Dastidar and Adeli (2009)  [15] introduced a novel supervised learning algorithm for multiple spiking neural networks, demonstrating their predictive capabilities in the clinically significant task of epilepsy and seizure detection. This study highlights the potential of SNNs for time-sensitive and pattern recognition tasks, laying groundwork for later applications in domains such as image classification and reservoir computing.
Nguyen et al. (2023)  [16] evaluated LIF-based spiking neural networks for image classification using the RAVSim simulator, demonstrating their predictive potential in structured visual tasks.
@@ 2.3 RC in Traffic Prediction
Further advancing the field, Zhang and Vargas (2023)  [18] applied reservoir computing in the context of 5G and beyond networks. Their study emphasized how reservoir computing architectures enhance forecasting accuracy while addressing sustainability challenges in mobile networks. By leveraging innovative learning techniques, their work provides actionable insights for integrating reservoir computing frameworks into modern telecommunications infrastructures, ensuring both predictive performance and energy efficiency. Xue et al. (2021)  [19] proposed a self-adaptive PSO-optimized Echo State Network for time series prediction, highlighting the potential of reservoir computing in modeling complex temporal patterns efficiently. Collectively, these studies underline reservoir computing’s transformative potential for cellular traffic forecasting, particularly in energy-constrained and dynamic environments like 5G networks. Its ability to balance computational efficiency with prediction accuracy makes it a critical tool for future network management and optimization.
@@ 2.4 Federated Bio Inspired in Traffic Prediction
Our study distinguishes itself by combining SNNs and RC models within federated learning environments to assess their predictive capacity and energy efficiency. Unlike prior studies that prioritize either centralized configurations or accuracy alone, we employ a Sustainability Index (See Section 3) to evaluate the trade-offs across multiple dimensions, including power consumption, prediction error, and data size. This comprehensive approach, coupled with comparative evaluations against traditional architectures such as MLPs and CNNs, provides deeper insights into how bioinspired models can support sustainable and scalable solutions for cellular traffic forecasting. By leveraging real-world datasets and consistent experimental setups, our work offers a holistic framework for advancing energy-efficient and privacy-preserving network management practices.
@@ 3.1 Dataset
As showed in [6], the dataset presents non Independent and Identically Distributed (non-IID) distibution, with the presence of three different skews, i.e. quantity, quality and temporality. This differentiation among respective Bias Sources (BS) is depicted in Fig. 3.1, indicating a challenging ML task.
@@ 3.3 ML Models
Reccurent Neural Network (RNN): RNNs are designed to capture temporal dependencies by maintaining a hidden state across sequential inputs. In our implementation, the RNN processes the time-series as an (S, d) sequence, where each step is passed through a recurrent layer with 128 hidden units, followed by a fully connected layer for prediction.
Lapicque Neuron (Original LIF Neuron): The Lapicque neuron implements the fundamental circuit approach, originating from biology. Ion channels of a physical neuron’s membrane are modeled with a resistor and the membrane itself is modeled with a capacitor. We check the potential ($U$) in the capacitor, and if this potential reaches a predefined threshold, the neuron fires an output spike. We also choose a mechanism that is activated when a spike is fired to reset our neuron. Thus, the tunable parameters in this approach are the resistance ($R$), capacitance ($C$), threshold, the number of spikes for each input ($I_{in}$) and the reset mechanism. Formally, Lapicque neuron [3] is modeled through the following function:
@@
When the resulting value of $U$ is higher than the specified threshold, the neuron will fire a spike. The threshold is a hyperparameter and it needs to be tuned to our problem to achieve, better predictive performance.
Leaky Neuron: The initial description of the LIF model, as described in [22], is distinguished by a broad set of hyperparameters. As the model is scaled to support a universal SNN architecture, managing these parameters becomes challenging. For this type of neuron, we use the following equation:
@@
Recurrent Leaky (RLeaky) Neuron: This implementation, as discussed in [3], employs a logic similar to that of Leaky neuron, yet it uniquely scales the output spikes of a neuron by a factor of $V$ and reincorporates them into the input. Specifically, spikes produced by a layer of spiking neurons are forwarded to the following fully connected (or convolutional) layer, where they are combined with the incoming input before being processed by the spiking neuron model.
Synaptic Based Neuron: In the neuron models discussed previously, the premise has been that a voltage spike upon input instantaneously elevates the synaptic current, subsequently influencing the membrane potential. However, in practical scenarios, a spike initiates a progressive discharge of neurotransmitters from the pre-synaptic to the post-synaptic neuron. The synaptic conductance-based [23] LIF model accommodates the slow temporal progression of the input current. Thus, the equation is transformed to:
@@
Reservoir Computing Echo State: Reservoir computing leverages a fixed, randomly initialized dynamic reservoir to project input data into a high-dimensional feature space. The Echo State Network (ESN) is a prominent implementation of this paradigm, particularly suited for sequential data processing. The reservoir consists of recurrently connected nodes, which transform the input time-series into a rich set of temporal features. Unlike traditional neural networks, the reservoir weights remain fixed, and only the output layer is trained, significantly reducing the computational complexity. In our implementation, the input data $(S,d)$ is mapped to the reservoir with a dimensionality of $N_{r}$ nodes. This dimensionality need to be tuned for our problem through a number of trials. The output is then computed as a linear combination of the reservoir states and passed to a fully connected layer for final predictions. This architecture facilitates efficient processing of multivariate time-series with minimal parameter tuning and low energy consumption. We selected Echo State Networks (ESNs) as a representative model of the reservoir computing paradigm due to their well-documented ability to handle temporal tasks with minimal training overhead. ESNs require training only the output layer, resulting in significantly reduced computational and energy demands—an important factor in our study of sustainable AI. Their simplicity and reproducibility also make them an ideal baseline for comparison. While other models within reservoir computing (e.g., Liquid State Machines) could be considered, our focus was to benchmark a widely adopted and computationally efficient architecture. In our experiments we use 128 reservoir neuron with fixed weights and only train the weights of the output layer.
@@ 3.4 Evaluation Metrics
The primary metric used to quantify prediction error for the time-series forecasting task is the normalized root mean squared error ($\mathrm{NRMSE}$), defined as $\mathrm{NRMSE}=\frac{1}{\overline{x}}\sqrt{\frac{\sum_{t=1}^{T}\left(\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}\right)}{Td^{\prime}}}$, where $\overline{x}$ represents the mean of the target observations, computed as $\overline{x}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\left(\sum_{i=1}^{d^{\prime}}x_{t,i}\right)$, and $\hat{x}_{t,i}$ and $x_{t,i}$ denote the predicted and actual values at time $t$ for feature $i$, $T$ is the total number of time steps, and $d^{\prime}$ is the number of features.
In addition to $\mathrm{NRMSE}$, we also employ other common error metrics to evaluate model performance: The Mean Absolute Error $\mathrm{MAE}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left|\hat{x}_{t,i}-x_{t,i}\right|$, the Root Mean Square Error $\mathrm{RMSE}=\sqrt{\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}}$ and the Mean Square Error $\mathrm{MSE}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}$.
Beside the accuracy of the implemented algorithms, we also utilize the Sustainability Indicator $\mathcal{S}$ introduced in [25] to assess the resulting trade-off between the prediction error and the energy consumption. $\mathcal{S}$ enables a fair comparison among different model architectures and can be calculated using the following formula:
@@
where $E_{\text{Val}}$ is the validation error (in this work, we consider NMRSE), $C_{\text{Tr}}$ represents the total energy consumed for model training in Watt Hours (Wh) and $D$ represents the size of the data exchanged during training. In centralized learning, $D$ corresponds to the total dataset size transferred to a central server. In contrast, in federated learning, $D$ denotes the size of the model weights exchanged between clients and the server, as only model updates are communicated rather than raw data. The use of this parameter, instead of direct power consumption during communication, is explained in [25]. This choice is primarily based on the fact that, although direct communication costs are often considered negligible  [26] , a broader sustainability analysis should also account for indirect factors—such as throughput and bandwidth requirements—which can impact the system’s overall energy footprint (e.g., increased cooling costs of data centers). The exponents denote the importance of each value, with $a+b+c=1$. The lower the value of $S$, the better the tradeoff between computational efficiency and predictive accuracy.
@@ 4 Experimental Setup
In the output layer, however, neurons lack a reset mechanism—a deliberate design choice to better address the requirements of the time-series forecasting task. This configuration ensures that the predicted output corresponds directly to the membrane potential, avoiding distortions caused by resets. Additionally, input spikes are encoded directly from the input data at specific timesteps, enabling the network to effectively convert non-spiking data into a format compatible with spiking mechanisms. This approach enhances the model’s ability to process time-series data within the spiking neural network framework. To ensure consistency, we employed the same dataset throughout all experiments. Moreover, we set $a=b=c=0.33$ to assign equal weight to power consumption, predictive performance, and size parameter in the index.
[…]
Best Hyperparameter Configuration for Each Model after Tuning via Optuna Python Library Model Hidden Neurons Threshold Learning Rate Batch Size Lapicque Neuron 256 1.6 $1.09\times 10^{-4}$ 130 Leaky Neuron 96 1.7 $2.39\times 10^{-4}$ 64 Recurrent Leaky Neuron 256 2.0 $5.85\times 10^{-4}$ 32 Synaptic Based Neuron 128 1.7 $1.07\times 10^{-4}$ 128 Alpha Neuron 256 2.0 $5.85\times 10^{-4}$ 32 Multi-Layer Perceptron 128 $10^{-3}$ 128 Convolutional Neural Network 128 $10^{-3}$ 128
For implementation, we utilized the PyTorch library [27] for traditional models such as MLPs and CNNs. The SNN models were implemented using the snnTorch library [3], which is tailored for gradient-based learning and provides pre-designed spiking neuron models that seamlessly integrate as recurrent activation units.
To ensure the validity and fairness of our comparative analysis, we employed Optuna Python Library  [28], that uses a Tree-structured Parzen Estimator (TPE) to tune key parameters in the centralized learning setup. Specifically, we optimized the number of hidden-layer neurons, threshold values, learning rate, batch size. The best-performing hyperparameter configurations identified through this process are summarized in Table  4. These configurations were subsequently applied to the federated learning setup to ensure consistency and comparability across different settings. We employed 150 epochs for the centralized learning, and 50 Federated Rounds with 3 Local epochs in the federated setup to ensure that the dataset was accessed the same number of times. Additionally, we incorporated an early stopping mechanism during training of both learning settings, halting the process if no improvement in the loss function was observed over 50 consecutive epochs. This not only prevented overfitting but also contributed to energy efficiency by reducing unnecessary computations.
Energy consumption during the training process was measured using the CarbonTracker Python module [29]. CarbonTracker is a tool designed to estimate and monitor the carbon emissions associated with computational processes, helping organizations and researchers assess and reduce their environmental impact. The experiments were conducted on a Windows 11 workstation equipped with an AMD Ryzen 7 6800HS CPU and 16 GB of memory.  [^2] The source code used in this paper is available at https://github.com/vperifan/Federated-Time-Series-Forecasting, accessed on May 10 2025.